# GeoLangBind: Unifying Earth Observation with Agglomerative Vision–Language Foundation Models

### v1

@@ 4.3 Modality-aware knowledge agglomeration
Modality-aware layer normalization
Given the feature map from the last layer of the GeoLangBind model, we first interpolate it to match the teacher’s feature resolution, denoted as $\mathbf{F}\in\mathbb{R}^{d\times H^{\prime}\times W^{\prime}}$, omitting the batch dimension for simplicity. With $\mathbf{F}$ and a modality prompt $\mathbf{V}_{p}\in\mathbb{R}^{d}$, our goal is to perform layer normalization that dynamically adapts to the modality. To achieve this, we introduce modality-aware conditional layer normalization, formulated as follows.
@@
Projector module
A convolutional layer is applied to adjust the feature representation $\mathbf{F}^{\prime}$ to channel dimensions $d_{t}$ of the corresponding teachers. The final representation is then used to compute the feature-matching loss for model distillation. Our design incorporates modality awareness into the feature translation process by encoding wavelength information and adapting the layer normalization parameters accordingly. This ensures that the learned features are both aligned and modality-specific, enhancing robust adaptation across different input modalities.
Model training
To train the model, we use a feature-matching loss to align the translated spatial features with the teacher’s features. As illustrated in Fig. 3, we extract three RGB channels from multispectral or hyperspectral images as input to the teacher model. The student features $\mathbf{F}_{s}\in\mathbb{R}^{d_{t}\times H^{\prime}\times W^{\prime}}$ are then matched to the teacher’s features $\mathbf{F}_{t}\in\mathbb{R}^{d_{t}\times H^{\prime}\times W^{\prime}}$ using a combination of $L_{1}$ loss, cosine embedding loss to maximize the cosine similarity, and mean squared error (MSE) loss:
We utilize three pre-trained models as teachers: SigLIP [64], DINOv2 [35], and ViT [13]. During training, the GeoLangBind model is initialized with SigLIP pre-trained weights and trained on GeoLangBind-2M using both the pairwise sigmoid loss $L_{\text{siglip}}$ for contrastive learning and distillation loss for feature matching. The total loss function is:
@@
[…]
Table 3: Zero-shot comparison of various models on GeoBench datasets. Bold values indicate the best performance.
@@
[…]
4.4 Progressive Multimodal Weight Merging
To address the data imbalance issue, we propose a post-training weight-space merging strategy to scale the model to diverse data modalities effectively. Since RGB data is the most dominant modality, we split the dataset into two parts: data with RGB channels and data with other modalities. During training, we first train two separate GeoLangBind models on these subsets. After training, we merge their weights to obtain a unified model capable of handling both RGB and other modalities. Since the original SigLIP model has been trained on massive data, we also merge the weights of GeoLangBind with the original SigLIP weights to enhance the model. Thus, we perform two rounds of weight merging to merge these three model weights: 1) merge the original SigLIP weights with the model trained on the RGB subset of GeoLangBind-2M (GeoLangBind-RGB); 2) merge the intermediate model with the model trained on the remaining five data modalities (GeoLangBind-Others).
For clarity, let $\theta_{\text{siglip}}$ be the original SigLIP model weights, $\theta_{\text{rgb}}$ be the GeoLangBind model weights trained on the RGB subset, $\theta_{\text{others}}$ be the GeoLangBind model weights trained on the other five modalities. In the first step, we merge $\theta_{\text{siglip}}$ and $\theta_{\text{rgb}}$ using a simple linear weight merging strategy:
$\theta^{*}=(1-m_{1})\,\theta_{\text{siglip}}+m_{1}\,\theta_{\text{rgb}},$ (7)
where $m_{1}$ controls the weighting ratio between the two models. Next, we merge the intermediate weights $\theta^{*}$ with $\theta_{\text{others}}$ to obtain the final GeoLangBind model:
$\theta=(1-m_{2})\,\theta^{*}+m_{2}\,\theta_{\text{others}},$ (8)
where $m_{2}$ determines the contribution of the non-RGB modalities. We conduct ablation studies in Sec. 5.3 by varying $m_{1}$ and $m_{2}$ to search for optimal weighting ratios. Although more sophisticated weight merging methods exist [3], we adopt a simple linear approach to evaluate its effectiveness for heterogeneous EO data. Investigating advanced merging strategies remains a topic for future work.
5 Experiments
5.1 Implementation details
GeoLangBind is trained using 8× NVIDIA A100 40GB GPUs for the base model and 8× NVIDIA A100 80GB GPUs for the large version. Our code is implemented based on OpenCLIP [21]. For the base version, we use a ViT-Base architecture (patch size 16), initialized with WebLI-pretrained weights [9]. The teacher models are ViT-Base-SigLIP (224), DINOv2-ViT-L/14, and ViT-Huge (ImageNet). The large version is based on SoViT-400m[4] (patch size 14, input size 384), with teacher models ViT-So400m-SigLIP (384), DINOv2-ViT-L/14, and ViT-Large (ImageNet). Training is conducted for 5 epochs on RGB datasets and 20 epochs on other modalities using AdamW (learning rate 5e-4, weight decay 1e-7). The batch size is 140 for the base model and 35 for the large model.
5.2 Performance comparison
We follow the evaluation protocols of SkyScript [50] and evaluate the models via zero-shot classification tasks. Specifically, the following eleven image scene classification datasets are used: AID [54], EuroSAT [19], fMoW [11], Million-AID [30], PatternNet [68], NWPU-RESISC45 [10], RSI-CB256 [25], as well as the fine-grained attribute classification datasets. In addition, we further compare these models on some datasets from the GEO-Bench suite [24]: m-bigearthnet [41], m-so2sat [71], and m-forestnet [22].
@@ 5.2 Performance comparison
AMFMN [61] 5.39 15.08 23.40 4.90 18.28 31.44 10.63 24.78 41.81 11.51 34.69 54.87 16.67 45.71 68.57 12.86 53.24 79.43
LW-MCR-u [60] 4.39 13.35 20.29 4.30 18.85 32.34 9.73 26.77 37.61 9.25 34.07 54.03 18.10 47.14 63.81 13.14 50.38 79.52
GaLR [62] 6.59 19.85 31.04 4.69 19.48 32.13 14.82 31.64 42.48 11.15 36.68 51.68
@@
[…]
Table 5: Partial fine-tuning results on six segmentation tasks. All models are trained with a frozen backbone for 20 epochs. Reported numbers are mean intersection over union (mIoU). Missing values are due to the inability of the model to adapt to this domain.
Method
Models
Backbone
m-pv4ger-seg
m-nz-cattle
m-NeonTree
m-cashew-plant
m-SA-crop
m-chesapeake
Supervised DeepLabv3 ResNet101 93.4 67.6 53.9 48.6 30.4 62.1
U-Net ResNet101 94.1 80.5 56.6 46.6 29.9 70.8
Self-supervised Vision FMs rand. init. ViT-B 81.7 74.1 51.7 32.4 29.0 47.1
Scale-MAE [38] ViT-L 83.5 76.5 51.0 - - 61.0
GFM [31] Swin-B 92.0 75.0 51.1 - - 63.8
Cross-Scale MAE [43] ViT-B 83.2 77.9 52.1 - - 52.3
CROMA [14] ViT-B - - - 30.1 31.4 -
DOFA [57] ViT-B 94.5 81.4 58.8 51.5 33.0 65.3
DOFA [57] ViT-L 95.0 81.8 59.4 56.9 32.1 66.3
Zero-shot FMs RemoteCLIP [28] ViT-L 94.1 79.8 56.2 37.9 31.3 59.5
Skyscript-20 [50] ViT-L 93.9 79.2 57.2 38.0 30.6 61.5
GeoLangBind-B-224 ViT-B 94.3 80.7 56.1 53.6 31.8 62.7
GeoLangBind-L-384 Large 94.7 82.2 59.0 45.8 32.6 63.0
@@
As shown in Tab. 2, our GeoLangBind models establish new state-of-the-art results in zero-shot remote-sensing image classification. Compared with existing CLIP-based baselines such as CLIP-original, RemoteCLIP, CLIP-laion-RS, and the various SkyCLIP configurations, GeoLangBind-B-224 (Base) consistently outperforms all counterparts on scene-level tasks. Its average accuracy across the eight scene datasets is significantly higher than that of the other Base models. The Larger variant of GeoLangBind with higher resolution yields further accuracy gains. This highlights that both capacity and resolution play vital roles in capturing the spatial and spectral details crucial for complex remote-sensing tasks. On three fine-grained roof-attribute classification tasks, both GeoLangBind variants achieve strong accuracy on roof shape and surface, significantly outperforming other methods. Although the results on smoothness remain more modest, our approach still provides a clear improvement over existing approaches.
Zero-shot classification on GEO-Bench
The zero-shot classification results on three GEO-Bench datasets are presented in Tab. 3. GeoLangBind demonstrates competitive or superior zero-shot multi-label classification performance compared to existing models on the m-bigearthnet dataset (43 classes). For the m-so2sat dataset (17 classes), when utilizing five spectral bands, the performance decreases slightly due to the significantly less multispectral training data compared to RGB. Note that competitive models cannot process 5-band images as input for zero-shot classification. On the m-forestnet dataset (12 classes), GeoLangBind substantially outperforms other models and shows remarkable accuracy improvements when using five spectral bands as input. Notably, the m-forestnet dataset uses Landsat-8 satellite data, which is not represented in our training dataset.
@@
Tab. 4 compares the GeoLangBind model against recent methods on three benchmark caption datasets: RSICD, RSITMD, and UCM-caption. We report Recall@1, Recall@5, and Recall@10 for both image-to-text and text-to-image retrieval. Despite the inherent difficulty of remote-sensing cross-modal tasks, GeoLangBind-L consistently achieves competitive or higher recall scores than most competitive baselines, including CLIP-original, CLIP-laion-RS, and SkyCLIP-30.
Semantic segmentation performance
Tab. 5 shows that GeoLangBind models work well across six segmentation tasks despite being zero-shot vision–language models. For all models, the backbone remains frozen while UPerNet [56] is trained as the segmentation head. GeoLangBind-B-224 achieves strong mIoU scores on m-pv4ger-seg and m-nz-cattle datasets. GeoLangBind-L-384 further improves results across all datasets, achieving competitive or better performance than vision foundation models. These results demonstrate GeoLangBind’s versatility and ability on fine-grained pixel-level understanding tasks.
@@
[…]
SigLIP+GeoLangBind-B-RGB 77.70 95.50
SigLIP+GeoLangBind-B-RGB 72.00 95.80
[…]
SigLIP+GeoLangBind-B-Others 69.50 96.20
SigLIP+GeoLangBind-B-Others 66.30 91.20
[…]
GeoLangBind-B-All-Merging 77.60 96.65
5.3 Ablation studies
@@
[…]
GeoLangBind-B w/ MaKA 62.3 82.2 59.0
Table 8: Ratio for merging SigLIP and GeoLangBind-B-RGB.
@@
Weight merging
Tab. 6 presents the results of our ablation experiments, evaluating the impact of weight merging across different models on the AID zero-shot classification dataset. The baseline SigLIP model can be significantly improved by incorporating the GeoLangBind-B-RGB model, trained on the RGB part of GeoLangBind-2M. Merging SigLIP with GeoLangBind-B-RGB further enhances performance. The model trained with our MaKA module (for distillation) can largely outperform the non-distilled version. This highlights the role of distillation in improving weight fusion. GeoLangBind-B-Others, trained on non-RGB modalities performs lower than its RGB counterpart. When merged with SigLIP, the performance can be significantly improved.
GeoLangBind-B-All is the model trained with all the data modalities mixing together, i.e. using data mix instead of weight merging. When compared with the GeoLangBind-B-All-Merging, the performance is significantly lower. This comparison demonstrates that weight merging is more effective and flexible than direct mixing data modalities.
Knowledge distillation
As presented in Tab. 7, we conducted ablation studies on the MADOS [23], m-nz-cattle [1], and m-NeonTree [52] datasets. GeoLangBind-B without MaKA represents a model trained using a baseline distillation method that does not incorporate modality-specific wavelengths as conditioning. The results demonstrate the significant effectiveness of the MaKA module, showing consistent performance improvements across all datasets.
@@
In Tab. 8 and Tab. 9, a linear search is conducted to determine the optimal ratios for weight merging. Based on the results, we select 0.9 for $m_{1}$ to merge the weights of SigLIP and GeoLangBind-RGB. We choose 0.5 for merging weights of the intermediate weights and GeoLangBind-Others. Thus we set $m_{1}=0.9$ and $m_{2}=0.5$ to derive our final model weights. The findings demonstrate that merged weights consistently outperform individual models (when the ratio equals 0 or 1), validating the effectiveness of our weight-merging approach.
Table 9: Ratio for merging weights of GeoLangBind-B-RGB and GeoLangBind-B-Others.
@@
Figure 4: t-SNE visualization of feature distributions for the original SigLIP features (left) and GeoLangBind-L-384 model (right). Circle markers represent language features, while triangle markers represent image features.
Figure 5: Visual comparison of deep features from RemoteCLIP, Skyscript, and GeoLangBind model.
5.4 Feature visualization and analysis
In the left side of Fig. 4, we present the t-SNE plot of SigLIP features. There is a significant separation between the language and image features for hyperspectral, SAR, elevation, and multispectral data modalities. This means that SigLIP struggles to align the representations of these data modalities. In contrast, features of GeoLangBind-L-384 demonstrate a much stronger feature alignment across all modalities, making it a powerful foundation for Earth observation tasks. In Fig. 5, we visually compare the feature heatmaps of RemoteCLIP, SkyScript, and GeoLangBind. We can see that GeoLangBind features are more spatially structured and preserve more object details.
6 Conclusion
We introduce GeoLangBind, a unified vision–language foundation model that bridges heterogeneous EO modalities through a shared language space. To achieve this, we construct GeoLangBind-2M, a large-scale image–text dataset containing six EO data modalities. GeoLangBind consists of a wavelength-aware encoder and a modality-aware knowledge agglomeration module to enhance fine-grained image understanding. Additionally, progressive weight merging is proposed to scale the model training to multiple data modalities. Extensive experiments demonstrate the state-of-the-art performance of GeoLangBind across zero-shot classification, semantic segmentation, and cross-modal retrieval tasks.
Appendix A Supplementary material

### v2

@@ 1. Wavelength-aware prompt generator
2. Modality-aware layer normalization
Given the feature map from the last layer of the DOFA-CLIP model, we first interpolate it to match the teacher’s feature resolution, denoted as $\mathbf{F}\in\mathbb{R}^{d\times H^{\prime}\times W^{\prime}}$, omitting the batch dimension for simplicity. With $\mathbf{F}$ and a modality prompt $\mathbf{V}_{p}\in\mathbb{R}^{d}$, our goal is to perform layer normalization that dynamically adapts to the modality. To achieve this, we introduce modality-aware conditional layer normalization as follows.
@@ 2. Modality-aware layer normalization
3. Projector module
A convolutional layer is applied to adjust the feature representation $\mathbf{F}^{\prime}$ to channel dimensions $d_{t}$ of the corresponding teachers. The final representation is then used to compute the feature-matching loss for model distillation. This design ensures that the learned features are both aligned and modality-specific, enhancing robust adaptation across different input modalities.
To address the data imbalance issue, we employ a weight space merging strategy (Ilharco et al. 2022). Since RGB data is the most dominant modality, we split the dataset into two parts: RGB data and non-RGB data. We first train two separate DOFA-CLIP models on these subsets. Then, we merge their weights to obtain a unified model that can handle both RGB and other modalities. Since the original SigLIP model is trained on massive data, we also merge the original SigLIP weights with our model. For clarity, let $\theta_{\text{siglip}}$ be the original SigLIP model weights, $\theta_{\text{rgb}}$ be the DOFA-CLIP model weights trained on the RGB subset, $\theta_{\text{others}}$ be the DOFA-CLIP model weights trained on the other five modalities. In the first step, we merge $\theta_{\text{siglip}}$ and $\theta_{\text{rgb}}$ using a simple linear weight merging strategy:
$\theta^{*}=(1-m_{1})\,\theta_{\text{siglip}}+m_{1}\,\theta_{\text{rgb}},$ (7)
where $m_{1}$ controls the weighting ratio between the two models. Next, we merge the intermediate weights $\theta^{*}$ with $\theta_{\text{others}}$ to obtain the final model:
$\theta=(1-m_{2})\,\theta^{*}+m_{2}\,\theta_{\text{others}},$ (8)
where $m_{2}$ determines the contribution of the non-RGB modalities. We conduct ablation studies by varying $m_{1}$ and $m_{2}$ to search for optimal weighting ratios. While more advanced merging methods exist (Akiba et al. 2025), we use a linear approach to assess its effectiveness on heterogeneous EO data, leaving sophisticated strategies for future work.
@@ 3. Projector module
DOFA-CLIP-L-384 76.83 75.50 59.04 29.10 70.16 80.17 73.15 51.62 64.45 61.83 26.00 81.36
Table 3: Zero-shot comparison on GeoBench datasets with non-RGB data. Bold values indicate the best performance.
@@
DOFA-CLIP-B (224) 47.70 20.37 23.69 17.95 14.60 13.60 17.02
Experiments
Implementation details
DOFA-CLIP is trained on 8 NVIDIA A100 40GB GPUs for the base model and 8 NVIDIA A100 80GB GPUs for the large version. For the base version, we use a ViT-Base architecture (patch size 16) pretrained on WebLI (Chen et al. 2022). The teacher models are ViT-Base-SigLIP (224), DINOv2-ViT-L/14, and ViT-Huge (ImageNet). The large version is based on SoViT-400m (Alabdulmohsin et al. 2023) (patch size 14, input size 384), with teacher models ViT-So400m-SigLIP (384), DINOv2-ViT-L/14, and ViT-Large (ImageNet). Training is conducted for 20 epochs using AdamW (learning rate 5e-4, weight decay 1e-7).
Performance comparison
We follow the evaluation protocols of Wang et al. (2024c) and evaluate the models via zero-shot classification tasks. Specifically, the following eleven image scene classification datasets are used: AID (Xia et al. 2017), EuroSAT (Helber et al. 2019), fMoW (Christie et al. 2018), Million-AID (Long et al. 2021), PatternNet (Zhou et al. 2018), NWPU-RESISC45 (Cheng, Han, and Lu 2017), RSI-CB256 (Li et al. 2017), as well as the fine-grained attribute classification datasets. In addition, we further compare these models on some datasets from the GEO-Bench suite (Lacoste et al. 2024): m-bigearthnet (Sumbul et al. 2019), m-so2sat (Zhu et al. 2019), and m-forestnet (Irvin et al. 2020).
@@
AMFMN (Yuan et al. 2022a) 5.39 15.08 23.40 4.90 18.28 31.44 10.63 24.78 41.81 11.51 34.69 54.87 16.67 45.71 68.57 12.86 53.24 79.43
LW-MCR-u (Yuan et al. 2021) 4.39 13.35 20.29 4.30 18.85 32.34 9.73 26.77 37.61 9.25 34.07 54.03 18.10 47.14 63.81 13.14 50.38 79.52
GaLR (Yuan et al. 2022b) 6.59 19.85 31.04 4.69 19.48 32.13 14.82 31.64 42.48 11.15 36.68 51.68
@@
DOFA-CLIP-L 8.42 25.16 37.05 8.15 24.90 37.73 13.94 30.31 44.47 13.45 38.70 55.78 43.33 81.43 93.81 35.28 71.62 87.80
@@
As shown in Table 2, DOFA-CLIP establishes new state-of-the-art results in zero-shot remote-sensing image classification. Compared with existing CLIP-based baselines such as CLIP-original, RemoteCLIP, CLIP-laion-RS, and the various SkyCLIP configurations, DOFA-CLIP-B-224 consistently outperforms all counterparts. Its average accuracy across the eight scene datasets is significantly higher than that of the other base models. The Larger variant of DOFA-CLIP with higher resolution yields further accuracy gains. On three fine-grained roof-attribute classification tasks, both DOFA-CLIP variants achieve strong accuracy on roof shape and surface, significantly outperforming other methods. Although the results on smoothness remain more modest, our approach still provides a clear improvement over existing approaches.
Zero-shot classification on multispectral imagery
The zero-shot classification results on three GEO-Bench datasets are presented in Table 3. DOFA-CLIP demonstrates competitive or superior zero-shot multi-label classification performance compared to existing models on the m-bigearthnet dataset (43 classes). For the m-so2sat dataset (17 classes), when utilizing five spectral bands, the performance decreases slightly due to the significantly less multispectral training data compared to RGB. Note that competitive models cannot process 5-band images as input for zero-shot classification. On the m-forestnet dataset (12 classes), DOFA-CLIP substantially outperforms other models and shows remarkable accuracy improvements when using five spectral bands as input. Notably, the m-forestnet dataset uses Landsat-8 data, which is unseen in our training dataset.
@@
Table 4 compares DOFA-CLIP against recent methods on three benchmark caption datasets: RSICD, RSITMD, and UCM-caption. We report Recall@1, Recall@5, and Recall@10 for both image-to-text and text-to-image retrieval. Despite the inherent difficulty of remote-sensing cross-modal tasks, DOFA-CLIP-L consistently achieves competitive or higher recall scores than most competitive baselines, including CLIP-original, CLIP-laion-RS, and SkyCLIP-30.
@@
DOFA-CLIP-B (RGB) 76.90 94.55
SigLIP+DOFA-CLIP-B (RGB) 77.70 95.50
SigLIP+DOFA-CLIP-B (RGB) 72.00 95.80
DOFA-CLIP-B (Others) 64.70 91.95
SigLIP+DOFA-CLIP-B (Others) 69.50 96.20
SigLIP+DOFA-CLIP-B (Others) 66.30 91.20
DOFA-CLIP-B (Mixed) (baseline) 71.60 93.65
DOFA-CLIP-B (Merging) (ours) 77.60 96.65
Ablation studies
@@
DOFA-CLIP-B w/o MaKA 61.7 81.5 58.1
DOFA-CLIP-B w/ MaKA 62.3 82.2 59.0
Table 7: Ratio for merging SigLIP and DOFA-CLIP-B (RGB).
@@
Weight merging and MaKA
Table 5 presents the results of our ablation experiments, evaluating the effects of weight merging and MaKA across different models on the AID zero-shot classification dataset. The baseline SigLIP model can be significantly improved by incorporating the DOFA-CLIP-B (RGB) model, trained on the RGB part of GeoLangBind-2M. The model trained with our MaKA module (for distillation) can largely outperform the non-distilled version. DOFA-CLIP-B (Others), trained on non-RGB modalities, performs lower than its RGB counterpart. When merged with SigLIP, the performance can be significantly improved. DOFA-CLIP-B (Mixed) is the model trained with all the data modalities mixed together. When compared with the DOFA-CLIP-B (Merging), the performance is significantly lower. This demonstrates that weight merging is more effective and flexible than direct mixing of data modalities.
Figure 4: t-SNE visualization of feature distributions for SigLIP features (left) and DOFA-CLIP-L-384 model (right).
MaKA enhances segmentation performance
As presented in Table 6, we conducted ablation studies on the MADOS (Kikaki et al. 2024), m-nz-cattle (Abuaiadah and Switzer 2022), and m-NeonTree (Weinstein et al. 2021) datasets. DOFA-CLIP-B without MaKA represents a model trained using a baseline distillation method that does not incorporate modality-specific wavelengths as conditioning. The results demonstrate the significant effectiveness of the MaKA module, showing consistent performance improvements across all datasets. More details are in the supplementary material.
@@
In Table 7 and 8, a linear search is conducted to determine the optimal ratios for weight merging. Based on the results, we select 0.9 for $m_{1}$ to merge the weights of SigLIP and DOFA-CLIP (RGB). We choose 0.5 for merging weights of the intermediate weights and DOFA-CLIP (Others). Thus, we set $m_{1}=0.9$ and $m_{2}=0.5$ to derive our final model weights. Note that the weight merging is executed a single time using a designated zero-shot classification dataset, and held fixed across all tasks.
Table 8: Ratio for merging weights of DOFA-CLIP-B (RGB) and DOFA-CLIP-B (Others).
@@
Figure 5: Visual comparison of deep features from RemoteCLIP, Skyscript, and DOFA-CLIP.
Feature visualization and analysis
On the left side of Fig. 4, we present the t-SNE plot of SigLIP features. Compared to the SigLIP features, the features of DOFA-CLIP-L-384 demonstrate a significantly stronger imagery-text alignment across all modalities, making it a powerful foundation for Earth observation tasks. In Fig. 5, we visually compare the feature heatmaps of RemoteCLIP, SkyScript, and DOFA-CLIP. DOFA-CLIP features are more spatially structured and preserve more object details, demonstrating the effectiveness of VECT.
Conclusion
We introduce DOFA-CLIP, a unified vision–language foundation model that bridges heterogeneous EO modalities through a shared language space. To achieve this, we construct GeoLangBind-2M, a large-scale image–text dataset containing six EO data modalities. DOFA-CLIP consists of a wavelength-aware encoder and a modality-aware knowledge agglomeration module to enhance fine-grained image understanding. Additionally, progressive weight merging is proposed to scale the model training to multiple data modalities. Extensive experiments demonstrate the state-of-the-art performance of DOFA-CLIP across zero-shot classification, semantic segmentation, and retrieval tasks.
Supplementary material